An electrician charges $15 plus $11 per hour. Another electrician charges $10 plus $15 per hour. Use a system of equations to find out for how many hours the electricians would charge the same amount.
Answer by mananth(16946)

An electrician charges $15 plus $11 per hour.
let number of hours be x
C= 15+11x

Another electrician charges $10 plus $15 per hour.
C= 10+15x
C should be same
so 15+11x=10+15x
4x=5
/4
x=5/4 days

Use a system of equations to find out for how many hours the electricians would charge the same amount.